The average first serve of a certain tennis player follows a symmetrical, mound-shaped distribution with a mean of 122 and a standard deviation of 6.2 (miles per hour). Based on this information, approximately 95% of the first serves will fall between 115.8 and 128.2.

True
False
Answer #1

Given that,

The average first serve of a certain tennis player follows a symmetrical, mound-shaped distribution with a mean of 122 and a standard deviation of 6.2 (miles per hour).

\mu-\sigma = 122-6.2 =115.8

\mu+\sigma = 122+ 6.2 =128.2

In normal distribution approximately 68% of the data values fall between 1 standard deviations from tje mean.

In above case, approximately 68% of the first serves will fall between 115.8 and 128.2

Therefore given statement is False